OpenStudy (anonymous):

Please Help! In the game of Greed, six dice are rolled. A player usually finds it desirable to roll at least one 5. (a) Find the probability at least one 5 occurs when all six dice are rolled. (b) When 5's occur, the player removes the dice showing a 5 and rolls again. If four 5's occur when all six dice are rolled, the remaining two are rolled. Find the probability of at least one 5 when the two dice are rolled.

OpenStudy (anonymous):

P(At least one 5)=1-P(No dies landed on a 5) We use the binomial distribution formula

OpenStudy (anonymous):

a) Pr(at least one 5) = 1 - Pr (no 5's) = 1 - (5/6)^6 = 0.6651 b) Pr(at least one 5) = 1 - Pr (no 5's) - 1 - (5/6)^2 = 0.3056

OpenStudy (anonymous):

Thank You! =)